0.78 g of sodium hydroxide (NaOH) are added to a 1 L solution of water at a temperature of 21°C in a tube separated by a semi-permeable membrane. Calculate the osmotic pressure of the solution.
1
Expert's answer
2020-04-22T09:17:31-0400

The osmotic pressure of a solution can be calculated according to the equation:

"\Pi = icRT" ,

using "i" - van't Hoff factor, "c" - the molar concentration of the solute, "R" - the ideal gas constant(0.082057 L·atm/mol·K), "T" - the temperature of the solution in kelvins.


Let's convert the temperature first:

"T = 21("°C") + 273.15 = 294.15" K.


The molar concentration of the solute (NaOH, molar mass is 40.00 g/mol) is:

"c = \frac{n}{V} = \frac{m}{M·V} = \frac{0.78(\text{g})}{40.00\text{(g/mol)}·1\text{(L)}} = 0.0195 \text{(mol/L)}".


The van't Hoff factor of strong electrolytes is equal to the number of ions produced from the dissociation of one formula unit. From NaOH formula unit, 2 ions are produced: Na+ and OH-. Therefore, the van't Hoff factor of NaOH is equal to 2.


Finally, the osmotic pressure is:

"\Pi = 2·0.0195 \text{(mol/L)}·0.082057 \text{(L atm/(mol K))}·294.15 \text{(K)} = 0.94 \text{ atm}"


Answer: the osmotic pressure of the solution is 0.94 atm.